### v2

# Who Connects Global Aid? The Hidden Geometry of 10 Million Transactions

Abstract
The global aid system functions as a complex and evolving ecosystem; yet widespread understanding of its structure remains largely limited to aggregate volume flows. Here we map the network topology of global aid using a dataset of unprecedented scale: over 10 million transaction records connecting 2,456 publishing organisations across 230 countries between 1967 and 2025. We apply bipartite projection and dimensionality reduction to reveal the geometry of the system and unveil hidden patterns. This exposes distinct functional clusters that are otherwise sparsely connected. We find that while governments and multilateral agencies provide the primary resources, a small set of knowledge brokers provide the critical connectivity. Universities and research foundations specifically act as essential bridges between disparate islands of implementers and funders. We identify a core solar system of 25 central actors who drive this connectivity including unanticipated brokers like J-PAL and the Hewlett Foundation. These findings demonstrate that influence in the aid ecosystem flows through structural connectivity as much as financial volume. Our results provide a new framework for donors to identify strategic partners that accelerate coordination and evidence diffusion across the global network.
Introduction
The modern architecture of global aid was born in the aftermath of World War II. In its infancy, Global Aid was conceived as a state-led instrument of geopolitical reconstruction managed by the OECD Development Assistance Committee[1]. It was driven primarily by diplomatic and domestic political imperatives, with the allocation often determined by donor strategic interests rather than recipient need [2, 3, 4]. However, the late 20th and early 21st centuries witnessed a radical phase transition[5]. A triple revolution characterised by new goals, new instruments and an explosion of new actors has reshaped the landscape[5]. This includes the rise of non-DAC donors[6], the proliferation of vertical funds and private foundations, and a shift toward the complex interdependence of the Sustainable Development Goals[7].
Despite the growing scale, the politics of measurement remain contentious[8]. High-level policy attempts to harmonise this sprawling system, such as the Paris Declaration[9], have largely failed. Empirical analysis suggests that coordination efforts have not reduced proliferation[10]. This leaves the sector in a state of unmanaged complexity, where uncoordinated high volumes of aid can sometimes hinder economic growth[11].
Today, the global aid system functions as a complex and evolving ecosystem. Our analysis of the International Aid Transparency Initiative (IATI) registry reveals a diverse population of 2,456 distinct publishing organisations ranging from multilateral banks and government agencies to private sector contractors and academic institutes. These actors operate across 230 countries and territories and utilise a complex mix of financial instruments including standard grants, aid loans and equity investments (Extended Data Fig. 1). The sheer scale of this system is now captured by over 10 million transaction records logged between 1967 and 2025. This complexity has outpaced the ability of policy makers to monitor the system through traditional statistical aggregates. In this environment outcomes emerge not from top-down directives but from millions of decentralised interactions between funders, implementers and intermediaries, that must be analysed as an autonomous, largely self-governed system.
Despite this structural complexity, the prevailing view of the aid sector remains one that ranks actors primarily by who spends the most. This perspective highlights the financial dominance of massive bilateral agencies and multilateral banks such as the World Bank or USAID. However, this volume-based view obscures the hidden geometry of the system. In a networked system, influence and efficiency are functions of connectivity and brokerage as much as financial weight. Focusing solely on volume masks the critical structural holes where coordination fails. It also hides the distinct islands of humanitarian and development activity that remain sparsely connected. To date, there has been no systemic map identifying the structural brokers responsible for bridging these functional and geographic clusters.
Here we present the first global-scale topological map of the aid system derived from a computational analysis of the IATI transaction corpus. By applying bipartite network projection and dimensionality reduction we reveal distinct functional clusters beyond traditional donor-recipient dyads. We identify a solar system architecture defined by the 100 most central actors with 25 at its very core. Crucially, we find that while governments and multilaterals dominate financial volume, a small set of knowledge brokers occupy disproportionately high bridging positions. We specifically identify universities and research foundations as the essential connective tissue holding the ecosystem together. We highlight the role of unanticipated central connectors such as J-PAL and the Hewlett Foundation. These findings suggest that to de-risk investments and accelerate evidence diffusion donors must distinguish between doers who deliver volume and connectors who maintain the structural integrity of the network.
Results
The Topology of Global Aid
To understand the physical footprint of the aid system we mapped the geographic distribution of over 10 million transaction records logged in the IATI registry between 1967 and 2025. Unlike traditional volume maps which highlight only a few massive recipients, the transaction density map (Figure 1) reveals a truly globalised system involving activity in 230 countries and territories. We observe a heavy long tail distribution. While a small number of nations like India and Bangladesh represent massive hubs of high-frequency activity the majority of the network consists of nations with moderate transaction volumes that are often invisible in aggregate financial reporting.
This geographic topology suggests a system that is widespread but highly fragmented. Activity is not evenly distributed but clustered around specific regional corridors. For instance we observe distinct high-density belts in East Africa and South Asia contrasted with sparser connectivity across Central Asia and South America. This fragmentation raises the question of whether these geographic clusters are integrated into a single cohesive network or function as separate islands of aid.
Recent network literature challenges the traditional volume hypothesis[12, 13, 14], suggesting instead that a recipient’s integration into the global network is a stronger predictor of growth and treaty ratification than raw financial flows[15, 16].
Figure 1: Global distribution of 10 million aid transactions from 1967 to 2025. Darker regions indicate higher transaction density on a logarithmic scale. This reveals the geographic footprint of the 2,456 publishing organisations in the International Aid Transparency Initiative (IATI) registry. While activity is global it is highly clustered around specific regional corridors in East Africa and South Asia with India and Bangladesh representing massive high-frequency hubs.
The Hidden Geometry of Funding and Implementation
To test for this structural integration, we applied Uniform Manifold Approximation and Projection (UMAP) to the network embeddings of all 2,456 provider organisations.
To derive these embeddings, we first constructed a bipartite graph consisting of nodes representing organisations of various types (e.g., Government, NGO, Private Sector) that either provide or receive fundings. The connections between these nodes are established by financial aid transactions, that is, specifically connecting organisations that provide funding to those that receive it. We then processed this bipartite graph structure using the node2vec algorithm. This technique converts the graph topology into high-dimensional embeddings (vectors) that capture the structural roles and connections of each organisation.
Finally, we used UMAP to project these high-dimensional vectors into a 2D space, revealing the hidden geometry of the sector. The analysis (Figure 2) exposes a system defined by two primary axes of differentiation rather than a cohesive whole.
Figure 2: The hidden geometry of the aid ecosystem. UMAP dimensionality reduction projects high-dimensional network embeddings into 2D space to reveal functional clusters defined by two primary axes of network behaviour. The horizontal axis represents a network tendency spectrum, separating actors with Humanitarian-like transaction patterns (left) from those with Development-like patterns (right). The vertical axis differentiates Funders (top) from Implementers (bottom). This orthogonal arrangement creates four distinct quadrants populated by exemplars such as OCHA (Humanitarian Funder), Save the Children (Humanitarian Implementer), IADB (Development Funder) and Chemonics (Development Implementer). The analysis also highlights the strategic position of connectors like J-PAL and The Hewlett Foundation, which sit near the origin and bridge these functional zones.
The horizontal axis reveals a persistent structural divide between Humanitarian and Development actors. Organisations involved in crisis response such as OCHA and Médecins Sans Frontières congregate on the left while long-term development finance institutions like the World Bank and USAID cluster on the right. This separation highlights the topological fissures that exist between short-term relief operations and long-term resilience building.
The vertical axis reveals a functional stratification between Funders and Implementers. Primary donors and government agencies occupy the upper hemisphere while implementing NGOs and private contractors populate the lower hemisphere. This orthogonal arrangement effectively divides the ecosystem into four distinct quadrants. For instance the upper-left quadrant contains humanitarian funders while the lower-right contains development implementers. While not every organisation conforms strictly to this schematic the pattern indicates that operational roles exert a strong influence on network topology.
The separation between these clusters highlights a systemic coordination risk. The structural holes between the quadrants are bridged by relatively few actors, potentially stalling the transition from crisis response to resilience building. We identify specific connectors such as J-PAL and The Hewlett Foundation that sit near the origin of this coordinate system, suggesting they are uniquely capable of integrating these otherwise disconnected functional zones. Similar topological fissures have been observed in global health networks[17] and climate adaptation finance[18], where issue-specific clusters often fail to integrate with broader systems strengthening[19, 20].
The Solar System of Brokerage
If the system is fragmented into islands, who acts as the bridge? We projected the network into a one-mode collaboration graph to identify the most central actors. We visualise the top 100 organisations (Figure 3) as a Solar System with actors ranked by their centrality in the network (see Extended Data Table LABEL:tab:tab1 for the full ranking).
To derive this ranking, we analysed the underlying bipartite graph connecting providers to receivers. We calculated the Network Centrality (Hub Score)[21] for all organisations to measure their systemic influence. In this bipartite context, a high Hub Score identifies actors that serve as critical “hubs”, connecting to a wide and significant range of recipients across the global network.
In the visualisation, nodes are sized to reflect the scale of organisations in terms of the total number of transactions, providing a visual contrast between activity volume (size) and structural influence (position). The “rings” are derived from the power-law distribution of these Hub Scores. The Inner Ring comprises the top 25 distinct “core” actors who exhibit disproportionately high structural influence compared to the long tail of the periphery.
We identify a core inner ring of 25 actors who dominate the network’s structural connectivity. This visualisation reveals a critical paradox between scale and influence. While the inner ring includes the expected financial giants (US Department of State, European Commission, World Bank) it also contains a distinct class of Knowledge Brokers. These are organisations with lower financial volume but disproportionately high network centrality (Extended Data Fig. 3).
Figure 3: The solar system architecture of global aid. Organisations are ranked by centrality (inner rings) and sized by transaction volume. While the core inner ring contains the expected financial giants (large nodes), it also reveals a distinct class of knowledge brokers (small nodes). Organisations such as J-PAL and The Hewlett Foundation occupy central positions despite having significantly smaller financial footprints than bilateral agencies. This highlights their strategic role as connectors rather than just funders.
Universities and Foundations as Bridges
Academic institutions are approximately 85% smaller by deal count than the global average (565 vs 3739), yet despite this, they occupy a significantly more central position with a median rank of 434 compared to the overall median of 571. This vital brokerage role is further evidenced by the fact that universities and research organizations constitute 81 of the top 1000 most central organisations and importantly 12 among the Top 100.
Foundations too, while less numerous than academic institutions in the global aid network with only 45 among the top 1000 and one in the Top 100, play a vital role as connectors. Qualitative analysis suggests these actors function as specialized innovation and knowledge brokers, bridging the structural gap between global financing and distinct technical or localized ecosystems. From setting data standards to scaling grassroots solutions, they provide essential connectivity across disparate thematic clusters.
We specifically highlight the most central foundation William and Flora Hewlett Foundation and academic institution: J-PAL (MIT) as critical connectors. Despite managing fewer deals than major bilaterals J-PAL’s network position allows it to diffuse evidence across the humanitarian-development divide more efficiently than many larger and siloed agencies.
The Hewlett Foundation’s pivotal role is visualized in its subnetwork (Extended Data Fig. 2), which reveals it supports a highly diverse portfolio of downstream partners. Our analysis confirms that Hewlett is directly connected to 12 of the top 100 most central organisations in the global network. These connections are significant and sustained, representing over 400 distinct transactions valued in the hundreds of millions. This portfolio bridges the gap between critical knowledge hubs like the Population Council (37 transactions) and J-PAL (70 transactions), and major implementation partners such as the United Nations Foundation (104 transactions) and Oxfam America (93 transactions).
Their high betweenness centrality [22] (Extended Data Fig.3) indicates both J-Pal and the Hewlett foundation sit on the shortest paths between otherwise poorly connected donor-implementer clusters. These findings overturn the assumption that influence is solely a function of budget. In the hidden geometry of global aid, smaller research-led organisations provide the essential connective tissue that holds the fragmented system together.
Discussion
Structural Fragmentation. Our topological analysis reveals that the global aid system is not a cohesive or integrated network. It is rather a collection of functional and geographic islands that are sparsely connected. The clear separation observed in the UMAP embeddings between the humanitarian and development clusters points to a persistent structural divide. Organisations tend to cluster with peers that share similar operational timeframes and funding sources[23, 24]. While this specialisation allows for operational efficiency within clusters it creates systemic risk at the macro level. It enables rapid response for humanitarian actors or long-term infrastructure planning for development banks but leaves the wider system disjointed.
The sparseness of connections between these clusters suggests a high potential for duplication of effort and coordination failure. The economic costs of such fragmentation are well documented, often overwhelming the bureaucratic capacity of recipient nations[25, 26] and creating game-theoretic equilibria where donors prioritise visibility over collective impact[27]. When evidence is generated in one cluster such as a randomised control trial on health interventions within the development cluster the lack of structural bridges prevents it from diffusing efficiently to the humanitarian cluster. This fragmentation explains a longstanding paradox in development economics regarding why evidence of what works often fails to scale. The network geometry confirms that the barriers to scaling are not merely financial or political[28, 24] but topological. Information simply cannot traverse the structural holes between the islands of the aid system without the intervention of specific bridging actors[29].
The Strategic Role of Brokers. The identification of a solar system architecture offers a new strategic framework for donors. Traditionally aid portfolios are constructed based on volume. Funding is allocated to doers such as large NGOs and contractors capable of absorbing significant capital. However our centrality analysis demonstrates that high-volume implementers often operate within deep but narrow silos. Industry analysis of donor practices confirms that this fragmentation creates a “vicious cycle” of grant dependence where short-term reporting requirements stifle the long-term systemic growth of support organisations[30].
To de-risk portfolios and accelerate systemic coordination donors must balance investments in doers with investments in connectors. Our results highlight that knowledge brokers provide disproportionate structural influence relative to their size. These are specifically academic research centers like the Abdul Latif Jameel Poverty Action Lab (J-PAL) at MIT and universities like Johns Hopkins University as well as strategic foundations like The Hewlett Foundation. These actors possess high betweenness centrality which means they sit on the shortest paths between otherwise disconnected nodes. Sociological theory posits that such weak ties are more effective for the diffusion of new information than strong ties to established partners[31]. This aligns with the rise of “knowledge actors” who exert influence through transnational standard-setting rather than financial weight[32, 33, 34]. J-PAL, for instance, leveraged this position to shift the sector toward evidence-based micro-interventions[35, 36, 37].
For donors this implies a shift in partnership strategy. When entering a new region or thematic area partnering with a high-volume implementer provides immediate operational capacity but limited reach. Conversely partnering with a central knowledge broker provides access to a pre-existing network of diverse downstream partners. By funding these brokers, donors essentially buy access to the structural integrity of the network. This allows for faster coordination and the de-risking of interventions through evidence-based consortia.
Network Resilience. From a complex systems perspective the current architecture of global aid exhibits signs of structural fragility. The system is heavily dependent on the Donor Core. This consists of a small number of massive hubs like USAID, the World Bank and the European Commission that provide the majority of financial fuel. In network theory systems dependent on a few central hubs are resilient to random failure but highly vulnerable to targeted disruption or the abrupt removal of a hub[38, 39]. This vulnerability is not hypothetical. Recent volatility such as the recalibration of US multilateral commitments between 2017 and 2021 demonstrates how policy discontinuities in a key donor nation can rapidly propagate instability across the wider network.
Strengthening the middle layer of knowledge brokers creates a more resilient and adaptive mesh. Much like the internet’s distributed architecture allows data to reroute around damaged nodes—unlike a centralised telephone exchange which fails if the switchboard goes down, a network with a strong middle layer of brokers is less prone to fragmentation even if a major donor pulls back. These brokers act as the system’s memory and nervous system by retaining evidence and maintaining relationships that transcend political cycles. Our analysis suggests that the resilience of the global aid architecture depends on shifting from a star topology of centralised control to a mesh topology of distributed coordination[40]. The hidden geometry we identified suggests that this mesh already exists in embryonic form through the university and foundation sectors. It simply requires recognition and resource support to function as a true stabiliser. Recent systemic design scholarship reinforces this arguing that traditional management dimensions must shift toward a “complexity paradigm” to handle such emergent dynamics[41].
Limitations. Our findings must be interpreted within the context of the data source. While IATI is the most comprehensive registry of global aid activity it is not exhaustive. Reporting is voluntary which leads to coverage heterogeneity. The dataset is biased toward organisations committed to transparency. Consequently flows from non-traditional donors, the private sector and certain non-reporting government agencies may be underrepresented. This suggests our map visualises the transparent aid network which may differ slightly from the total aid network.
Furthermore network links in this study represent financial transactions or reported collaborations. These are proxies for relationships but they do not guarantee the quality of the partnership. A single transaction may represent a one-off payment rather than a sustained collaboration. Additionally temporal lags in reporting mean the network map reflects a trailing indicator of activity rather than a real-time pulse. Similarly, the functional quadrant mapping is not prescriptive or exhaustive; while some organisations may appear in quadrants distinct from their nominal function due to hybrid operations, the analysis reveals a clear structural and statistical pattern in the ecosystem at large. Finally the definition of brokerage relies on betweenness centrality. This is a structural metric that assumes information flows along the shortest paths which may not always capture the complex political reality of aid negotiation.
Conclusion. We are moving from an era of Aid as Charity defined by the unidirectional transfer of volume to an era of Aid as Network defined by multi-directional connectivity. This study provides the first empirical evidence that the physical volume of aid does not equate to structural influence. By mapping the 10 million transactions that comprise the modern aid system we have revealed a hidden geometry where influence is wielded by those who connect rather than just those who spend.
The discovery of a core ecosystem of knowledge brokers offers a viable path forward for a fragmented sector. If the goal of global development is to solve complex cross-border challenges from climate resilience to pandemic preparedness the system must evolve from a collection of isolated islands into a cohesive whole. This requires a fundamental revaluation of the connector. In the complex ecosystem of global aid the organisations that weave the network together are just as critical as those that supply the capital.
Methods
Data Source and Processing
We derived the global aid transaction network from the International Aid Transparency Initiative (IATI) registry[20], the most comprehensive open-data standard for development cooperation reporting. Transaction-level data were programmatically retrieved using the IATI API, covering the full temporal span of available records from 1967 to 2025.
The data collection pipeline iteratively queried the IATI transactions endpoint with pagination to ensure complete coverage across all reporting organisations. Each API response was parsed and stored in a structured format, retaining core fields including transaction identifiers, reporting organisation, provider organisation, receiver organisation, transaction date, transaction value, currency, and transaction type. To ensure consistency across reporting practices, all transaction values were filtered to exclude null or zero-value entries, and only valid financial flows were retained for analysis. Organisation names in the raw IATI records exhibit substantial heterogeneity due to inconsistent reporting conventions, abbreviations, and spelling variants. To address this, we applied a systematic normalisation procedure combining exact matching and fuzzy string matching to resolve common aliases (e.g., mapping “USAID” to “United States Agency for International Development”). This process reduced duplication and ensured that each organisation was represented as a unique node in the network.
Transactions were then aggregated at the organisation–organisation level, linking provider and receiver entities through observed financial flows. Records with missing or unresolvable organisation identifiers were excluded. The final cleaned dataset comprised over 10 million transactions involving 2,456 distinct publishing organisations operating across 230 countries and territories. All preprocessing steps were implemented in Python to ensure reproducibility. The resulting cleaned and normalised transaction table served as the basis for subsequent network construction, projection, and analysis.
Network Construction
We modelled the global aid ecosystem as a bipartite (two-mode) network derived from the cleaned IATI transaction dataset. In this representation, nodes correspond to organisations, and edges represent observed financial transactions between a provider organisation and a receiver organisation[42, 43, 44]. Each transaction induces a directed edge from provider to receiver, reflecting the flow of resources within the aid system.
To construct the network, we first grouped transactions by unique provider–receiver organisation pairs. Multiple transactions between the same pair were aggregated into a single weighted edge, with the edge weight corresponding to transaction frequency rather than monetary value. This choice emphasises sustained relational activity and repeated collaboration over raw financial volume, allowing the network topology to capture structural connectivity rather than scale alone.
The resulting bipartite graph includes all organisations that appear at least once as either a provider or a receiver in the transaction corpus. Self-loops and records with unresolved organisation identifiers were excluded. The graph was implemented using the NetworkX library, enabling efficient handling of large-scale relational data and the computation of downstream structural metrics.
Dimensionality Reduction and Clustering
To reveal latent structural patterns in the global aid network, we applied a two-stage representation-learning and dimensionality-reduction pipeline based on node embeddings, followed by a nonlinear projection. This approach enables the comparison of organisations based on their topological roles rather than their geographic location or financial volume.
We first generated high-dimensional vector representations of organisations using the node2vec algorithm[45] applied to the bipartite graph. Node2vec performs biased random walks over the network to learn embeddings that capture both local neighbourhood structure and broader structural equivalence. Walks were generated from each node and optimised using a skip-gram objective, producing fixed-length embeddings for all organisations in the network. These embeddings encode similarities in funding and implementation patterns, allowing organisations with comparable network positions to be mapped close together in embedding space, even if they are not directly connected. The resulting high-dimensional embeddings were then projected into two dimensions using Uniform Manifold Approximation and Projection (UMAP)[46]. UMAP is a nonlinear dimensionality reduction technique that preserves local neighbourhood relationships while maintaining meaningful global structure. We used a cosine distance metric on the embedding vectors, with standard parameter settings (including a neighbourhood size of 15 and a minimum distance of 0.1), to obtain a stable two-dimensional manifold representation of the network.
The 2D projection reveals a clear geometric organisation of the aid ecosystem, in which organisations cluster by functional role and patterns of interaction. Rather than imposing predefined categories, clusters emerge organically from similarities in network connectivity. These clusters correspond to systematic differences between humanitarian and development actors, as well as between funding-oriented and implementation-oriented organisations. This embedding-based approach enables the identification of structural proximity and separation within the aid system without reliance on exogenous labels. The resulting low-dimensional representation forms the basis for the visual analyses presented in the Results section. It supports the identification of structural bridges and coordination gaps across the global aid network.
Centrality and Brokerage Analysis
We calculated network centrality metrics to identify the core architecture of the system. Degree Centrality was used to measure the total connectivity of actors. Betweenness Centrality was calculated to quantify brokerage[47]. This identifies actors that lie on the shortest paths between other pairs of organisations[48]. We validated these offline network metrics by correlating them with online visibility metrics using network centrality measures of linkage patterns from the organisations’ websites. We found a strong positive Pearson correlation of 0.48 between the structural position of an organisation in the IATI network and the centrality of its website in the World Wide Web (Global PageRank) (Extended Data Fig. 4). The solar system visualisation ranks the top 100 organisations by a composite centrality score with the inner ring representing the top 25 actors. To handle the scale of the graph we employed advanced approximation algorithms for neighbourhood functions[49]. This methodological approach aligns with previous efforts to map aid flows[50, 51, 52] but extends the analysis to a global scale, allowing for the detection of critical transitions[53] and systemic risks[54, 55].
Data Availability
The dataset analysed during the current study is publicly available in the International Aid Transparency Initiative (IATI) Registry. The raw XML data can be accessed at IATI API https://api.iatistandard.org/datastore. The processed network files and edge lists used for the topological analysis are available in the IATI_Global_Aid repository at https://github.com/leagueofscholars/IATI_Global_Aid.git.
Acknowledgements
We thank the International Aid Transparency Initiative (IATI) for maintaining and providing open access to the global aid data that made this research possible. We also acknowledge Snowmelt for its support and for valuable discussions that helped inform the broader context of this work.
Author contributions statement
P.X.M. conceived the study, developed the core ideas, and led the writing of the manuscript. X.G. collected and curated the data and contributed to the empirical analysis. M.-A.R. and P.B. provided conceptual guidance, methodological insights, and critical feedback that shaped the analysis and interpretation. All authors reviewed and approved the final manuscript.
Additional Information
The author(s) declare no competing interests.
Supplementary Information
Extended Data Fig. 1: Evolution of aid instruments (1967–2025). Longitudinal analysis of deal counts and instrument types (grants, loans, equity) demonstrating the increasing complexity and scale of the system over five decades.
Extended Data Fig. 2: The Hewlett Foundation Aid Graph (US\$1.7Bn, n=3,494). The diagram illustrates the Foundation’s support of 289 diverse entities, highlighting a strategic alignment with highly central actors in the global aid ecosystem. The network includes twelve top-100 ranked organizations, such as GlobalGiving, J-PAL, and the International Rescue Committee, bridging the gap between broad-based support and high-impact central nodes.
Extended Data Fig. 3: Comparative betweenness centrality. This chart contrasts the betweenness centrality scores of J-PAL and The Hewlett Foundation against the median of the top 25 most central organizations. The significantly higher scores for these knowledge brokers identify them as positive outliers that bridge disconnected clusters more effectively than typical central nodes.
Extended Data Fig. 4: Correlations between the Aid Graph Hub Score and other centrality measures. The chart displays Pearson correlation coefficients comparing the Hub Score against internal graph metrics (blue) and external global web metrics (green). The high correlation with internal PageRank (r=0.95) validates the metric’s consistency within the aid network. The moderate correlation with Global Website PageRank (r=0.48) indicates a measurable link between an organization’s “offline” co-investment centrality and its “online” authority on the World Wide Web.
Extended Data Table 1: Top 100 Organisations. Full tabular list of the Top 100 organisations ranked by composite centrality, including organisation type and sector.
Rank Organisation Name Website Organisation type
1 GlobalGiving.org globalgiving.org International NGO
2 The William and Flora Hewlett Foundation hewlett.org Foundation
3 IntraHealth International intrahealth.org International NGO
4 Abdul Latif Jameel Poverty Action Lab at MIT povertyactionlab.org Academic, Training and Research
5 International Initiative for Impact Evaluation 3ieimpact.org International NGO
6 Innovations for Poverty Action (IPA) poverty-action.org International NGO
7 UK: Foreign, Commonwealth and Devt Office gov.uk Government
8 United States Health and Human Services $h$s.gov Government
9 Ireland: Department of Foreign Affairs dfa.ie Government
10 United States Department of State state.gov Government
11 MSI Reproductive Choices msichoices.org International NGO
12 Guttmacher Institute guttmacher.org International NGO
13 World Health Organization (WHO) who.int Multilateral
14 Denmark: Ministry of Foreign Affairs, Danida um.dk Government
15 The Population Council, Inc. popcouncil.org Academic, Training and Research
16 International Labour Organization (ILO) ilo.org Multilateral
17 International Organization for Migration (IOM) iom.int Multilateral
18 Catholic Relief Services crs.org International NGO
19 UN Office for Project Services (UNOPS) unops.org Multilateral
20 UN Foundation unfoundation.org National NGO
21 Johns Hopkins University jhu.edu Academic, Training and Research
22 UN Development Programme (UNDP) undp.org Multilateral
23 PwC pwc.co.uk Private Sector
24 International Planned Parenthood Federation ippf.org International NGO
25 DT Global UK dt-global.com Private Sector
26 UN Joint Programme on HIV and AIDS unaids.org Multilateral
27 London School of Hygiene and Tropical Medicine lshtm.ac.uk Academic, Training and Research
28 ZOA zoa-international.com International NGO
29 UN Women unwomen.org Multilateral
30 ODI Global odi.org Academic, Training and Research
31 World Resources Institute wri.org International NGO
32 African Development Bank afdb.org Multilateral
33 Asian Development Bank adb.org Multilateral
34 UN Office Coord of Humanitarian Affairs unocha.org Multilateral
35 Malaria Consortium malariaconsortium.org International NGO
36 Liverpool School of Tropical Medicine lstmed.ac.uk Academic, Training and Research
37 GOAL goalglobal.org International NGO
38 Intl Institute for Sustainable Development iisd.org International NGO
39 PATH path.org International NGO
40 European Commission: International Partnerships ec.europa.eu Other Public Sector
41 Gavi: the vaccine alliance gavi.org Public Private Partnership
42 CARE International UK careinternational.org.uk International NGO
43 The International Rescue Committee rescue.org International NGO
44 Ipas ipas.org International NGO
45 Cardno Emerging Markets cardno.com Private Sector
46 ICF Consulting Services Ltd icf.com Private Sector
47 MannionDaniels manniondaniels.com Private Sector
48 Danish Refugee Council drc.ngo International NGO
49 UK: Dept of Health and Social Care (DHSC) gov.uk Government
50 Institute of Development Studies ids.ac.uk Academic, Training and Research
51 International Rescue Committee UK rescue-uk.org International NGO
52 Sightsavers sightsavers.org International NGO
53 British Council britishcouncil.org Other Public Sector
54 Plan International UK plan-uk.org International NGO
55 Clinton Health Access Initiative clintonhealthaccess.org International NGO
56 Health Poverty Action healthpovertyaction.org International NGO
57 Options Consultancy Services options.co.uk Private Sector
58 Intl Institute for Environment and Development iied.org Academic, Training and Research
59 International Medical Corps UK internationalmedicalcorps.org.uk International NGO
60 Restless Development restlessdevelopment.org International NGO
61 Leonard Cheshire Disability leonardcheshire.org International NGO
62 The HALO Trust halotrust.org International NGO
63 Saferworld saferworld.org.uk International NGO
64 Farm Africa farmafrica.org International NGO
65 World Agroforestry Centre worldagroforestry.org Academic, Training and Research
66 The World Bank worldbank.org Multilateral
67 TechnoServe technoserve.org International NGO
68 Intl Development Law Organization (IDLO) idlo.int Multilateral
69 International Potato Center cipotato.org Multilateral
70 Itad itad.com Private Sector
71 Camfed International camfed.org International NGO
72 BRAC brac.net International NGO
73 The Asia Foundation asiafoundation.org International NGO
74 Zoological Society of London zsl.org International NGO
75 Cadmus International UK Limited nathanlondon.co.uk Private Sector
76 Norwegian People’s Aid npaid.org International NGO
77 Hivos hivos.nl International NGO
78 Pact pactworld.org International NGO
79 Elrha elrha.org International NGO
80 CANADEM canadem.ca International NGO
81 Global Witness globalwitness.org International NGO
82 University of Cape Town uct.ac.za Academic, Training and Research
83 Stockholm International Water Institute siwi.org International NGO
84 HealthNet TPO healthnettpo.org International NGO
85 AECOM aecom.com Private Sector
86 CLASP clasp.ngo International NGO
87 Makerere University mak.ac.ug Academic, Training and Research
88 Ethiopian Public Health Association etpha.org Other Public Sector
89 Crown Agents Limited crownagents.com Private Sector
90 Girls’ Education Challenge: Fund Manager PwC girlseducationchallenge.org Private Sector
91 Save the Children UK savethechildren.org.uk International NGO
92 Mercy Corps Europe europe.mercycorps.org International NGO
93 Cordaid cordaid.org International NGO
94 The University of Manchester manchester.ac.uk Academic, Training and Research
95 Mott MacDonald Limited mottmac.com Private Sector
96 Disasters Emergency Committee dec.org.uk National NGO
97 Concern Worldwide UK concern.org.uk International NGO
98 Action Against Hunger UK actionagainsthunger.org.uk International NGO
99 Social Development Direct Limited sddirect.org.uk Private Sector
100 Voluntary Service Overseas (VSO) vsointernational.org International NGO